How old is a sample of wood whose 14C content is found to be 43% that of a living tree? The half-life of 14C is 5730 years.


(b) Iodine-131 is one of the most important isotopes used in the diagnosis of thyroid cancer. One atom has a mass of 130.906114 amu. Calculate the binding energy per nucleon. 


1
Expert's answer
2021-04-08T12:09:32-0400

Decay constant  "=\frac{0.693}{t\frac{1}{2}}= \frac{0.693}{5730years}=
1.209×10^{-4}/years"


The rate of counts is proportional to the number of C-14 atoms in the sample.

"N_0\u200b=100,"

"N=43"

The age of the sample "=t=\\frac{2.303}{k}\u200blog[\\frac{N_0}{N}]"


"t=\\frac{2.303}{1.209\u00d710^{-4}}\u00d7\u200blog[\\frac{100}{43}]=" 4430 years





Binding energy per nucleon. 


"=\frac{1.184481amu.\frac{931Mev}{1amu}}{131}=8.422Mev/nucleons"